how do you find the point of intersection of two straight-lines if there are no numbers that ccan cancel each other out. Example: y = −6x − 8; y = 7x + 3
Answer by richwmiller(17219)

in this case set the two equal to each other and solve for x
−6x − 8=7x + 3
x = -11/13,
and plug that in to get y
y = -38/13